Given that vector A =5i-7j+3k and B=-4i+4j-8k,find (I)AxB(ii)BxA
1
Expert's answer
2020-12-06T17:16:43-0500

Given,

"A =5\hat{i}-7\hat{j}+3\hat{k},"


"B=-4\hat{i}+4\hat{j}-8\hat{k}"


"A\times B = \begin{bmatrix}
\hat{i} & \hat{j}& \hat{k}\\
5 & -7 & 3\\
-4 & 4 & -8\\
\end{bmatrix}"


"=\hat{i}(56-12)-\hat{j}(-40+12)+\hat{k}(20-28)"


"=44\hat{i}+28\hat{j}-8\hat{k}"


"B\times A = \begin{bmatrix}
\hat{i} & \hat{j}& \hat{k}\\
-4 & 4 & -8\\
5 & -7 & 3\\
\end{bmatrix}"


"=\hat{i}(12-56)-\hat{j}(-12+40)+\hat{k}(28-20)"


"=-44\hat{i}-28\hat{j}+8\hat{k}"